OpenStudy (iloveleatiforever):

Solve for p. 2p + 2 = 8

OpenStudy (leenathan):

p = 3 2*3=6 then 6+2=8

OpenStudy (iloveleatiforever):

u sure this time

OpenStudy (leenathan):

yes

OpenStudy (iloveleatiforever):

ok thnx<3

OpenStudy (veronicachance):

Cuz 8-2=6 and 6/2= 3

OpenStudy (iloveleatiforever):

thnx